OpenStudy (anonymous):

solve the logarithmic equation: ln x +ln(x+1)=5

hartnn (hartnn):

hi @chelle1018 :) do you know the log property that \(\ln A+\ln B = \ln (AB)\) ?

hartnn (hartnn):

using that property , left side will equal \(\ln [x(x+1)]\) got this ?

OpenStudy (anonymous):

so: \[e^5=x(x+1)\] or \[e^5=x^2+x\]

OpenStudy (anonymous):

actually you can rewrite it as:\[\frac{ e^5 }{ x }-x=1\]

OpenStudy (anonymous):

so:\[\frac{ e^5 }{ x }-x=e^0\]

hartnn (hartnn):

x^2+x -e^5 =0 is quadratic equation, solve it using quadratic formula

OpenStudy (anonymous):

multiply both sides by x:\[e^5-x^2=x\]

OpenStudy (anonymous):

if you plug that quadratic equation into the quadratic formula you get \[x=\frac{ -1-\sqrt{1+4e^5} }{ 2 }\]

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

or \[x=\frac{ \sqrt{1+4e^5}-1 }{ 2 }\]